OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 2 3 All real numbers No solution

hero (hero):

First step...divide everything by -2 to get: x + 3 = x + 1 + 2 then simplify

OpenStudy (anonymous):

need help

OpenStudy (mathstudent55):

Since you have a number, -2 outside parentheses, you can multiply it out using the distributive property.

hero (hero):

When you say you "need help", do you mean that you "need the answer"? @supanick14

OpenStudy (anonymous):

no how to work it out

OpenStudy (mathstudent55):

The distributive property is: a(x + y) = ax + ay When a number or a variable multiplies a sum, multiply the nunmber or variable by each term of the sum.

OpenStudy (mathstudent55):

On the left side, you have -2(x + 3) = -2*x + (-2) * 3 = -2x - 6 Now do the same to the -2 and the parentheses on the right side.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

its two

OpenStudy (mathstudent55):

It isn't. Now you need to also multiply on the right side: −2(x + 1) − 4 = = -2x -2 - 4 = -2x - 6 So you end up with the equation: -2x - 6 = -2x - 6

hero (hero):

@supanick14, you should show the work you did to get the two.

OpenStudy (anonymous):

ok -2(x+1)-4 -2x-2-4 +2 +2 -2x-6=-2x=-4 x-6=-4 +6 +6 x=2

OpenStudy (mathstudent55):

Let's start from the beginning with the whole equation at once. −2(x + 3) = −2(x + 1) − 4 -2x - 6 = -2x - 2 - 4 -2x - 6 = -2x - 6 Now continue. What do you get?